GoConqr has a new look! Let us know what you think.

Flashcards

Probability

ella rose Maths English 9 cards

Cards in this deck (9)

  1. 1PROBABILITY Number of ways an outcome can happen Total number of possible outcomesIf tossing a coin, what is probability of getting heads?
    Number of ways you can get heads = 1
    Total possible outcomes = heads and tails = 2
    Probability = 1/2 = 50%
    When you throw a fair dice there are six possible outcomes: 1, 2, 3, 4, 5 or 6. There are three ways of getting an odd number (1, 3 or 5).
    So the probability of getting an odd number equals the number of ways of getting an odd number, divided by the total number of possible outcomes:
    Equation: = 3/6 = 1/2 or 50%
  2. 2modemost common average
    add all factors together and divide by number of factors
  3. 3mediantake the middle of the factors
  4. 4modethe most common of the various factors
  5. 5mutually exclusivemeans that two events cannot both be true at the same time
  6. 6not mutually exclusivemeans that both events CAN be true at the same time
  7. 7Example: I have a bag of letters spelling the word MOOSE. If i pick 1 letter out, what is the probability of it being: (a) O (b) a vowel (c) a TANSWER:
    (a) 2/5 - very unlikely
    (b) 3/5- likely
    (c) 0/5- impossible
  8. 8Theoretical possibilitywhen you estimate on the basis of a theory
  9. 9Experimental probabilitywhen you do an experiment to TEST your theory

Information

Description
Maths Flashcards on Probability, created by ella rose on 18/05/2015.
Created by
ella rose
Views
1
Rating

Feedback

What is it about?

What do you think of the new look?

How do you like it?